OpenStudy (anonymous):

A given spring has a spring constant of 400 newtons per meter. How much work is done in stretching the spring a distance of 0.02 meter from its equilibrium position?

OpenStudy (anonymous):

\[work =1\div 2 timesk \times e \times e\]

OpenStudy (anonymous):

therefore\[1/2 \times 400 \times 0.2\times 0.2\]

OpenStudy (anonymous):

solve that,you gonna get it

OpenStudy (anonymous):

why equation? isn't the equation for work is W=Fd?

OpenStudy (anonymous):

no..thats if you are given force and distance

OpenStudy (anonymous):

but you are given a spring constand so d formula is W=1/2 k e square

OpenStudy (anonymous):

and one more question, how much work is done to stretch the spring an additional 0.02 meter?

OpenStudy (anonymous):

den add 0.2 to 0.2..dat wil giv u 0.4

OpenStudy (anonymous):

so multiply 0.4 insted of that 0.2 in the first place

OpenStudy (anonymous):

thank you so much!!